OpenStudy (anonymous):

whats are the horizontal and vertical asymptotes for y=3^x+3 ?

OpenStudy (tkhunny):

Do you mean \(y = 3^{x} + 3\), as you have written it? Or, did you mean something else?

OpenStudy (anonymous):

its as i've written it, just like in my book

OpenStudy (anonymous):

the exponent is x+3

OpenStudy (tkhunny):

Then it is NOT as you have written it. You should have written y = 3^(x+3). This is the same as \(y = 3^{x+3}\). Do you see the difference it makes when adding the parentheses? You must remember your Order of Operations.

OpenStudy (anonymous):

yes, i see that. thank you. sorry

OpenStudy (tkhunny):

No worries. Writing in-line is very different from formatted printing in a book. Think very hard on your Order of Operations. Make sure you write what you mean. \(y = 3^{x+3}\) What is the LEAST value that can be obtained for y? Can you find a value for x that gives y = 1? How about a value for x that gives y = 0?

OpenStudy (anonymous):

x has to =0 to get y=1

OpenStudy (tkhunny):

Okay, we managed y = 1. How about y = 0? Can we do that?

OpenStudy (anonymous):

cant get it

OpenStudy (anonymous):

the answer says that y>0. but i cant see how they get that

OpenStudy (tkhunny):

How close can you get? \(y = 3^{-3+3} = 3^{0} = 1\) -- We got that, already. \(y = 3^{-4+3} = 3^{-1} = 1/3\) \(y = 3^{-5+3} = 3^{-2} = 1/9\) \(y = 3^{-6+3} = 3^{-3} = 1/27\) Do you see where this is going? Will it EVER get negative? How close to zero can we make it?

OpenStudy (anonymous):

oh. so the restriction is to always use positive values. otherwise we cant get anywhere. thank you so much!

OpenStudy (tkhunny):

Not USE positive values for y, but observe that you can't GET a negative value for y. You can snuggle up to y = 0, as close as you like, but you can't quite get there!

OpenStudy (anonymous):

yeah, i see that now. thanks :)